OpenStudy (anonymous):

Find the equation of the linear function f using the given info: f(0)=-3, f(3)=-3 I got a slope of -3, plugged in and got y = x - 3. Is this correct?

OpenStudy (anonymous):

How did you get slope = -3?

OpenStudy (anonymous):

Somebody told me "since both y values are the same, evidently it is a constant, namely f(x)=−3" and to plug that in

OpenStudy (anonymous):

But that's not how you find the slope. You find the slope by using the slope formula: (y2-y1/x2-x1) Do you see where y2, y1, x2, and x1 are? Those are your two sets of coordinates. Use them to find the slope, then you can plug it into the y = mx + b and find b, which will give you everything you need for your equation.

OpenStudy (anonymous):

OH, so y = 2x + b and now put in x and y from either set of points

OpenStudy (anonymous):

I plugged in 0,0... got 0 = 2(0) + b

OpenStudy (anonymous):

What slope did you get? Show me your calculations for slope.

OpenStudy (anonymous):

4 - 0 = 4 -2 - 0 = -2 -4/2 = -2/1 which equals -2. My bad

OpenStudy (anonymous):

I still get b = 0 regardless, so y = -2x + 0?

OpenStudy (anonymous):

Where did you get 4 and 2 from? Your functions are f(0)=-3 and f(3)=-3.

OpenStudy (anonymous):

So in the first one, x = 0, y = -3. And in the second one x = 3, y = -3. Those are your y2, x2, y1, x1.

OpenStudy (anonymous):

Oh crap I'm sorry, was looking down at my book at the question right below the one I'm going over. it was f(0)=0, f(-2)=4. Well, got that one right now lol.

OpenStudy (anonymous):

Doing the right one now

OpenStudy (anonymous):

-3 + 3 = 0 3 - 0 = 3 0/3?

OpenStudy (anonymous):

So what is your slope? Any fraction with 0 as the numerator equals what?

OpenStudy (anonymous):

undefined?

OpenStudy (anonymous):

No, that's if there's a 0 in the denominator. 0/3 is a number. A 0 in the numerator of any fraction just makes the fraction equal to 0.